OpenStudy (anonymous):

On Monday Andrew sold 15 packages of donuts. On Tuesday he sold 28 packages of donuts. What is the percent of increase, to the nearest percent, in donut sales from Monday to Tuesday? A. 87% B. 0.8% C. 86% D. .877%

OpenStudy (anonymous):

28-15 / 15 = 13/15 = .866666666 multiply by 100 to get percent = 86.666666 round to nearest percent = 87%

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Thank you both so much!

OpenStudy (anonymous):

ur welcome